# von Kármán–Howarth Similarity of Spatial Correlations and the Distribution of Correlation Lengths in Solar Photospheric Turbulence

### v1

@@ Abstract
Fluctuations in the Sun’s photospheric magnetic field are the primary source of the turbulence that can heat and accelerate the solar atmosphere, and thus play an important role in the production and evolution of the solar wind that permeates the heliosphere. A key parameter that characterizes this turbulence is the correlation scale of fluctuations, which determines the injection of turbulent energy into the plasma and influences the diffusive transport of solar energetic particles. This study employs magnetogram data acquired by the Helioseismic and Magnetic Imager on board the Solar Dynamics Observatory to characterize an ensemble of spatial autocorrelation functions (ACFs) of magnetic fluctuations in the photosphere. It is shown that the two-point ACFs satisfy the similarity-decay hypothesis of von Kármán and Howarth, a fundamental property of turbulent systems: following a rescaling of the ACFs by energy and correlation lengths, a quasi-universal functional form is obtained demonstrating exponential decay of correlations. The probability distribution function of transverse correlation lengths ($\lambda$) is shown to be approximately log-normal, which is consistent with observations of magnetic turbulence in the solar wind. A “mosaic” of the spatial distribution of $\lambda$ over the photosphere is presented; the “quiet Sun” tends to have $\lambda\sim 1500$ km (albeit with a wide distribution), which is close to the scale of solar granulation; systematically longer lengths are associated with active regions. A positive correlation is observed between mean magnetic field magnitude and $\lambda$, and empirical fits are derived to quantify this relationship. These results improve our understanding of the nature of turbulence in the solar photosphere and the origin of coronal and solar-wind turbulence, while reinforcing the universality of fundamental properties of turbulent flows observed in the natural world. The results also provide observational constraints for models that describe the transport of turbulence from solar and stellar photospheres into their coronae and beyond.
Keywords: Sun: photosphere – Sun: magnetic fields – turbulence
@@ 2 Theoretical and Observational Background
Like most properties of a turbulent system, the probability distribution function (PDF) of correlation lengths tends to show a large variance. Observational studies of this PDF in interplanetary space suggest a lognormal form, i.e., $\log(\lambda)$ has a normal distribution (Ruiz et al., 2014; Isaacs et al., 2015; Pradata et al., 2025). Lognormal distributions are widely observed in the physical, biological, and social sciences; they are associated with systems containing a large number of independent multiplicative (non-linear) effects (in contrast to the additive effects that produce normal distributions), leading to asymmetric PDFs with a long tail towards large values higher than the mean (e.g., Limpert et al., 2001). The lognormality of correlation scales has been linked to the generation of “1/f1/f” noise in the low-frequency ($f$) spectrum of the heliospheric magnetic field (Wang et al., 2024a, and references within).
@@ 4 Results
III. Probability Distribution of Correlation Lengths. Fig. 3 shows a histogram (top panel) and a PDF (bottom panel) of the correlation lengths computed above [using Eq. (2)]. Here $\lambda$ includes both $\lambda_{x}$ and $\lambda_{y}$ to obtain a larger statistical sample; their separate PDFs (not shown) are quite similar. The histogram, which is shown with a linear scale on the horizontal axis, contains the telltale one-sided tail towards large values of $\lambda$, a suggestion of a lognormal PDF (Limpert et al., 2001). The most probable value is $\sim 1500$ km. The PDF plot is shown for $\log\lambda$ on the bottom axis, with selected values of $\lambda$ displayed on the top axis. Values of $\lambda$ above 5000 km are excluded due to insufficient counts (see top panel). The plot contains two PDFs shown as a “stacked” bar plot: blue bars with thick outlines show the PDF after excluding $\lambda$ values associated with ARs, while the orange region above or below the top of each blue bar indicates the modification to the PDF when ARs are included in the ensemble. Note that for $\log\lambda\lesssim 7.5$ ARs tend to “pull down” the PDF while above this value they tend to increase the PDF, indicating that large values of $\lambda$ are associated with ARs, as also seen in Fig. 2. This is consistent with the recent study of ARs by Abramenko & Suleymanova (2024).
A best-fit of the three-parameter Gaussian function $A_{0}\exp{[(X-A_{1})^{2}/(2A_{2}^{2})]}$ to the blue PDF (ARs removed) is computed, where $X\equiv\log\lambda$. A nonlinear least-squares fit yields $A_{0}=1.6,~A_{1}=7.1~\text{and}~A_{2}=0.2$. [^5] The IDL function gaussfit.pro is used. $A_{0},A_{1},$ and $A_{2}$ are the height, center, and standard deviation of the Gaussian, respectively, with $1\sigma$ error estimates 0.09, 0.02, and 0.02. This function is plotted as a dashed-red curve, while the green curve shows an “arithmetic” Gaussian, i.e., one with the same mean and standard deviation as the blue PDF. The figure indicates that the PDF of correlation lengths (excluding ARs) is approximately described by a lognormal distribution. The corresponding fit to the PDF that includes ARs is less good (not shown), suggesting that turbulent fluctuations in the quiet Sun may constitute a different population from those in active regions. A more detailed and systematic study of the goodness of fits of the PDF (e.g., Ruiz et al., 2014) will require a larger sample (utilizing a large number of full-disk magnetograms covering an extended time period), and will be taken on in future work. We note here that the PDFs of sizes of solar supergranules have also been found to be lognormal (Noori et al., 2019).
@@
IV. Correlation of $\lambda$ with Mean Magnetic Field Strength. We have seen that magnetic fluctuations in ARs have larger correlation lengths than the QS, as has also been noted in other recent work (Abramenko & Suleymanova, 2024). This could potentially indicate differences in turbulence dynamics occuring in QS and ARs. At the same time, since ARs have larger magnetic field magnitudes than the QS, our results also suggest a correlation between the large-scale magnetic field $B$ and $\lambda$. To probe this further we plot in Fig. 4 a scatter plot of $\lambda$ (combining $\hat{\bm{x}}$ and $\hat{\bm{y}}$ components) as a function of the mean magnitude of the magnetic field $\langle|B|\rangle$, where averaging is performed over the same domains as those for the respective ACFs (Sec. 4.I). A significant correlation is apparent, with a computed Pearson correlation coefficient (e.g., Hoel, 1960) of 0.77. QS and ARs are represented with distinct symbols, reinforcing the previous findings of large $\lambda$ in the latter. Even in the QS (blue circles), a positive correlation between $\lambda$ and magnetic field is evident.
@@ 5 Conclusions and Discussion
(length scales are in km) Photosphere Chromosphere & T.R. Coronal base Young solar wind Near-Earth
$1~R_{\odot}$ 1.001 – 1.01 $R_{\odot}$ 1.01 – 1.3 $R_{\odot}$ 0.1 – 0.6 AU 1 AU 2 – 40 AU
$\perp$ Mag. corr. length 1500 (a) 1500 – 5000 (e) 3500 – 8000 (f) $10^{4}~\text{--}~0.7\times 10^{6}$ (g) $10^{6}$ (h) $10^{6}\text{--}10^{7}$ (i)
$\perp$ Mag. corr. length in a coronal hole 1000 (b)
$\perp$ Mag. corr. length in ARs $8000~\text{--}~40,000$ (c)
Granulation scale 500–2000 (d)
Supergranulation scale 30,000 (d)
Table 1: Various length scales and their evolution with distance from the Sun. All lengths are in km and are approximate values; see sources (below) for details on uncertainties and statistical spreads. Here $\perp$ Mag. corr.” stands for perpendicular (or transverse) magnetic correlation, defined relative to a mean magnetic field direction, and T.R. stands for transition region. Note that the top row refers to a general correlation length, not necessarily associated with a particular type of photospheric structure. Sources of values in the table are as follows. (a) Chhiber et al. (2025, present work). (b) Abramenko et al. (2013). (c) Abramenko & Suleymanova (2024); Chhiber et al. (2025, present work). (d) Rincon & Rieutord (2018), and references within. (e) Bailey et al. (2025). (f) Sharma & Morton (2023); Bailey et al. (2025); Hahn et al. (2025). (g) Ruiz et al. (2014); Chhiber et al. (2021b); Cuesta et al. (2022). (h) Ruiz et al. (2014); Cuesta et al. (2022). (i) Smith et al. (2001); Breech et al. (2008); Ruiz et al. (2014); Adhikari et al. (2017); Cuesta et al. (2022).
@@
III. We confirmed the validity of the von Kármán–Howarth similarity hypothesis for photospheric magnetic turbulence. This hypothesis is the basis for phenomenological treatments of turbulence decay of the form
@@
IV. Our results offer new insights on the origin and evolution of turbulence in the heliosphere. Table 1 lists approximate values of the correlation length at different distances from the Sun, starting from $\sim 1500$ km at the photosphere, increasing to $\sim 5000\text{--}8000$ km at the coronal base, $\sim 10^{4}\text{--}10^{6}$ through the young solar wind up to 1 AU, and reaching $\sim 10^{7}$ km by 40 AU. This increase is likely a combined effect of geometrical expansion effects (e.g., Hollweg, 1986) and in-situ development of turbulence (e.g., Matthaeus et al., 1998; Bruno & Carbone, 2013; DeForest et al., 2016; Chhiber et al., 2018). The approximately lognormal PDF of correlation lengths in the photosphere suggests that non-linear multiplicative processes are occurring at the source of the solar wind, with further evolution of the PDF via in-situ processes taking it closer to a lognormal form as the solar wind expands into the inner heliosphere (see Ruiz et al., 2014). These results have implications for our understanding of “$1/f$ noise” in frequency spectra across the heliosphere (e.g., Wang et al., 2024a). In particular a well-known generic path to producing $1/f$ spectra is the superposition of powerlaw signals with a scale-invariant weighting. But it has also been shown that a lognormal distribution can closely approximate a scale-invariant distribution over a number of decades in frequency (Montroll & Shlesinger, 1982). Therefore the presence of lognormality in photospheric correlation scales supports the idea that $1/f$ signals might originate in the solar dynamo (Wang et al., 2024a, and references within). We note that the von Kármán–Howarth similarity collapse of correlation functions in the photosphere also supports the notion that the solar photosphere is in a turbulent state (Petrovay, 2001), which in turn reflects on possible turbulent properties in the solar interior (Rincon & Rieutord, 2018).
@@ Appendix B Correlation Length in relation to the Magnetic Energy Spectrum
Figure 6: Blue curve shows power spectral density of line-of-sight magnetic field on the photosphere as a function of wavenumber $k$, computed as described in the text. Red dashed and green dash-dotted lines indicate reference $k^{-5/3}$ and $k^{-1}$ spectra, respectively. The inertial-range $\sim-5/3^{\text{rd}}$ spectrum transitions into the shallower “1/k1/k” range at scales slightly larger than the approximate correlation length on the quiet Sun, with the latter indicated with a vertical grey line at $k=1/2000~\text{km}^{-1}$.

### v2

@@ Abstract
Fluctuations in the Sun’s photospheric magnetic field are the primary source of the turbulence that can heat and accelerate the solar atmosphere, and thus play an important role in the production and evolution of the solar wind that permeates the heliosphere. A key parameter that characterizes this turbulence is the correlation scale of fluctuations, which determines the injection of turbulent energy into the plasma and the diffusive transport of solar energetic particles. This study employs magnetogram data from the Helioseismic and Magnetic Imager on the Solar Dynamics Observatory to characterize an ensemble of spatial autocorrelation functions (ACFs) of turbulence in the photosphere. It is shown that the two-point ACFs satisfy the similarity-decay hypothesis of von Kármán and Howarth, a fundamental property of turbulent systems: rescaling the ACFs by their respective energies and correlation lengths yields a quasi-universal exponential form. The probability distribution function of transverse correlation lengths ($\lambda$) is shown to be approximately log-normal, which is consistent with observations of turbulence in the solar wind. A “mosaic” of the spatial distribution of $\lambda$ over the photosphere is presented; the “quiet Sun” tends to have $\lambda\sim 1500$ km (albeit with a wide distribution), which is close to the scale of solar granulation; systematically longer lengths are associated with active regions. A positive correlation is observed between mean magnetic field magnitude and $\lambda$, and empirical fits quantify this relationship. These results improve our understanding of solar turbulence while providing observational constraints for models that describe turbulence transport from solar and stellar photospheres into their atmospheres.
keywords:
Sun: photosphere – Sun: magnetic fields – turbulence
pubyear: 2025pagerange: von Kármán–Howarth Similarity of Spatial Correlations and the Distribution of Correlation Lengths in Solar Photospheric Turbulence8
@@ 2 Theoretical and Observational Background
Like most properties of a turbulent system, the probability density function (PDF) of correlation lengths tends to show a large variance. Observational studies of this PDF in interplanetary space suggest a lognormal form, i.e., $\log(\lambda)$ has a normal distribution (Ruiz et al., 2014; Isaacs et al., 2015; Pradata et al., 2025). Lognormal distributions are widely observed in the physical, biological, and social sciences; they are associated with systems containing a large number of independent multiplicative (non-linear) effects (in contrast to the additive effects that produce normal distributions), leading to asymmetric PDFs with a long tail towards large values higher than the mean (e.g., Limpert et al., 2001). The lognormality of correlation scales has been linked to the generation of “1/f1/f” noise in the low-frequency ($f$) spectrum of the heliospheric magnetic field (Wang et al., 2024a, and references within).
@@ 4 Results
III. Probability Distribution of Correlation Lengths. Fig. 3 shows a histogram (top panel) and a PDF (bottom panel) of the correlation lengths computed above [using Eq. (2)]. Here $\lambda$ includes both $\lambda_{x}$ and $\lambda_{y}$ to obtain a larger statistical sample; their separate PDFs (not shown) are quite similar. The histogram, which is shown with a linear scale on the horizontal axis, contains the telltale one-sided tail towards large values of $\lambda$, a suggestion of a lognormal PDF (Limpert et al., 2001). The most probable value is $\sim 1500$ km. The PDF plot is shown for $\log\lambda$ on the bottom axis, with selected values of $\lambda$ displayed on the top axis. Values of $\lambda$ above 5000 km are excluded from the PDF due to insufficient counts (see top panel). The plot contains two PDFs shown as a “stacked” bar plot: blue bars with thick outlines show the PDF after excluding $\lambda$ values associated with ARs, while the orange region above or below the top of each blue bar indicates the modification to the PDF when ARs are included in the ensemble. Note that for $\log\lambda\lesssim 7.5$ ARs tend to “pull down” the PDF while above this value they tend to increase the PDF, indicating that large values of $\lambda$ are associated with ARs, as also seen in Fig. 2. This is consistent with the recent study of ARs by Abramenko & Suleymanova (2024).
A best-fit of the three-parameter Gaussian function $A_{0}\exp{[-(X-A_{1})^{2}/(2A_{2}^{2})]}$ to the blue PDF (with ARs removed) is computed, where $X\equiv\log\lambda$. This nonlinear least-squares fit yields $A_{0}=1.6,~A_{1}=7.1~\text{and}~A_{2}=0.2$. [^5] The IDL function gaussfit.pro is used. $A_{0},A_{1},$ and $A_{2}$ are the height, center, and standard deviation of the Gaussian, respectively, with $1\sigma$ error estimates 0.09, 0.02, and 0.02. This function is plotted as a dashed-red curve, while the green curve shows an “arithmetic” Gaussian, i.e., one with the same mean and standard deviation as the blue PDF. The figure indicates that the PDF of correlation lengths (excluding ARs) is approximately described by a lognormal distribution. The corresponding fit to the PDF that includes ARs is less good (not shown), suggesting that turbulent fluctuations in the quiet Sun may constitute a different population from those in active regions. A more detailed and systematic study of the goodness of fits of the PDF (e.g., Ruiz et al., 2014) will require a larger sample, utilizing a large number of full-disk magnetograms that cover an extended time period, and will be taken on in future work. We note here that the lognormal distribution has been shown to describe other elements of the photosphere, including the sizes of supergranules (Noori et al., 2019) and distributions of magnetic flux (Abramenko & Longcope, 2005).
Following the characterization scheme suggested by Limpert et al. (2001) for lognormal distributions (which they call multiplicative normal distributions), we report the geometric mean of $\lambda$:
@@
We also report the mean and standard deviation of the lognormally distributed $\lambda$ in terms of the parameters $A_{1}$ and $A_{2}$ of the best-fit Gaussian computed above (Aitchison & Brown, 1957; Abramenko & Longcope, 2005): $\text{mean}~(\lambda)\equiv\exp\left(A_{1}+A_{2}^{2}/2\right)=1237$ km and $\text{standard deviation}~(\lambda)\equiv\exp(2A_{1}+2A_{2}^{2})-\exp(2A_{1}+A_{2}^{2})=250$ km.
IV. Correlation of $\lambda$ with Mean Magnetic Field Strength. We have seen that magnetic fluctuations in ARs have larger correlation lengths than the QS, as has also been noted in other recent work (Abramenko & Suleymanova, 2024). This could potentially indicate differences in turbulence dynamics occuring in QS and ARs. At the same time, since ARs have larger magnetic field magnitudes than the QS, our results also suggest a correlation between the large-scale magnetic field $B$ and $\lambda$. To probe this further we show in Fig. 4 a scatter plot of $\lambda$ (combining $\hat{\bm{x}}$ and $\hat{\bm{y}}$ components) as a function of the mean magnitude of the magnetic field $\langle|B|\rangle$, where averaging is performed over the same domains as those for the respective ACFs (Sec. 4.I). A significant correlation is apparent, with a computed Pearson correlation coefficient (e.g., Hoel, 1960) of 0.77. QS and ARs are represented with distinct symbols, reinforcing the previous findings of large $\lambda$ in the latter. Even in the QS (blue circles), a positive correlation between $\lambda$ and magnetic field is evident.
@@ 5 Conclusions and Discussion
($\perp$ mag. corr. len. Photosphere Chromosphere & T.R. Coronal base Young solar wind Near-Earth
in km) $1~R_{\odot}$ 1.001 – 1.01 $R_{\odot}$ 1.01 – 1.3 $R_{\odot}$ 0.1 – 0.6 AU 1 AU 2 – 40 AU
Corr. length 1500 (a) 1500 – 5000 (d) 3500 – 8000 (e) $10^{4}~\text{--}~0.7\times 10^{6}$ (f) $10^{6}$ (g) $10^{6}\text{--}10^{7}$ (h)
Corr. length in a coronal hole 1000 (b)
Corr. length in ARs $8000~\text{--}~40,000$ (c)
Table 1: $\perp$ correlation lengths of magnetic turbulence in the photosphere, and evolution with distance from the Sun. For comparison, the length scales associated with granules and supergranules are 500-2000 km and 30,000 km, respectively (Rincon & Rieutord, 2018, and references within). All lengths are in km and are approximate values; see sources (below) for details on uncertainties and statistical spreads. Here $\perp$ mag. corr. len.” stands for perpendicular magnetic correlation length, defined relative to a mean magnetic field direction, and T.R. stands for transition region. Note that the top row refers to a general correlation length, not necessarily associated with a particular type of photospheric structure. Sources of values in the table are as follows. (a) Chhiber et al. (2025, present work). (b) Abramenko et al. (2013). (c) Abramenko & Suleymanova (2024); Chhiber et al. (2025, present work). (d) Bailey et al. (2025). (e) Sharma & Morton (2023); Bailey et al. (2025); Hahn et al. (2025). (f) Ruiz et al. (2014); Chhiber et al. (2021b); Cuesta et al. (2022). (g) Ruiz et al. (2014); Cuesta et al. (2022). (h) Smith et al. (2001); Breech et al. (2008); Ruiz et al. (2014); Adhikari et al. (2017); Cuesta et al. (2022).
@@
III. We confirmed the validity of the von Kármán–Howarth similarity hypothesis for photospheric magnetic turbulence. The correlation functions of highly dissimilar regions in the photosphere (see also Abramenko et al., 2013; Abramenko & Suleymanova, 2024) collapse to a quasi-universal exponential form following a rescaling by their respective energies and correlation lengths. The exponential nature of the correlation function itself has deep roots in terms of diagnosing the state of self-organization of a system (Watkins et al., 2016). We note, however, that while the exponential decorrelation describes the large-scale end of the inertial range and the energy-containing (and larger) scales, the smaller-scale behavior (near the dissipation scale) may be different (Taylor, 1938; Bandyopadhyay et al., 2020). Of course, probing the small-scale end of the inertial range in a reliable manner is challenging with current magnetogram resolutions.
The vK-H hypothesis is the basis for phenomenological treatments of turbulence decay of the form
@@
IV. Our results offer new insights on the origin and evolution of turbulence in the heliosphere. The vK-H similarity collapse supports the notion that the solar photosphere is in a turbulent state (Petrovay, 2001), which in turn reflects on possible turbulent properties in the solar interior (Rincon & Rieutord, 2018). Table 1 lists approximate values of the correlation length at different distances from the Sun, starting from $\sim 1500$ km at the photosphere, increasing to $\sim 5000\text{--}8000$ km at the coronal base, $\sim 10^{4}\text{--}10^{6}$ through the young solar wind up to 1 AU, and reaching $\sim 10^{7}$ km by 40 AU. This increase is likely a combined effect of geometrical expansion effects (e.g., Hollweg, 1986) and further in-situ development of turbulence (e.g., Matthaeus et al., 1998; Bruno & Carbone, 2013; DeForest et al., 2016; Chhiber et al., 2018). Note that the top row of Table 1 refers to a general correlation length, not necessarily associated with a particular type of photospheric structure. Future studies that focus on the radial evolution of correlation lengths associated with specific photospheric sources may “fill out” the second and third rows of the Table.
The approximately lognormal PDF of correlation lengths in the photosphere suggests that non-linear multiplicative processes are occurring at the source of the solar wind, with further evolution of the PDF via in-situ processes taking it closer to a lognormal form as the solar wind expands into the inner heliosphere (see Ruiz et al., 2014). These results have implications for our understanding of “$1/f$ noise” in frequency spectra across the heliosphere (e.g., Wang et al., 2024a). In particular a well-known generic path to producing $1/f$ spectra is the superposition of powerlaw signals with a scale-invariant weighting. But it has also been shown that a lognormal distribution can closely approximate a scale-invariant distribution over a number of decades in frequency (Montroll & Shlesinger, 1982). Therefore the presence of lognormality in photospheric correlation scales supports the idea that $1/f$ signals might originate in the solar dynamo (Wang et al., 2024a, and references within).
@@ Appendix B Correlation Length in relation to the Magnetic Energy Spectrum
Figure 6: Blue curve shows power spectral density of line-of-sight magnetic field in the photosphere as a function of wavenumber $k$, computed as described in the text. Red dashed and green dash-dotted lines indicate reference $k^{-5/3}$ and $k^{-1}$ spectra, respectively. The inertial-range $\sim-5/3^{\text{rd}}$ spectrum transitions into the shallower “1/k1/k” range at scales slightly larger than the approximate correlation length on the quiet Sun, with the latter indicated with a vertical grey line at $k=1/2000~\text{km}^{-1}$.